# revtex4-2Repair the float

### v1

@@ III Models
To model this hierarchical structure, we will use the LogisticPCA [16] and the new MSM. We selected LPCA as the representative for the SSM whereas, at the best of our knowledge, only the MSM accounts for a renormalization rule over arbitrary partitions of the nodes.
@@ III.1 Logistic PCA
LPCA [^2] The authors have historically called LPCA the directed LPCA, but there was no straightforward acronym for its undirected counterpart [16]. Since in this work, we will use only the symmetric one, we will call it LPCA for easiness. [16] is a probabilistic model which aims to classify each edge as existing (class $1$) or non-existing (class 00). This is a common “spam / non-spam” problem [13], but applied to objects (the links) that are not carrying explicit features - that is why node embeddings come into play.
The connection probability reads
@@
were introduced by the Chanpuriya et colleagues to grasp the “homophily” and “heterophily” natures of a node respectively (see Supp.Mat. I). By defining the non-negative matrices as $\textbf{B}:=\begin{bmatrix}-\vec{b}_{1}^{T}-\\ \vdots\\ -\vec{b}_{N_{0}}^{T}-\end{bmatrix}$ and similarly for $\mathbf{C}$, the scalar product among them becomes $\mathbf{X}:=\mathbf{B}\mathbf{B}^{T}-\mathbf{C}\mathbf{C}^{T}$ implying $x_{ij}:=\langle\vec{b}_{i},\vec{b}_{j}\rangle-\langle\vec{c}_{i},\vec{c}_{j}\rangle\in\mathbb{R}$. Therefore, the node embeddings are obtained by maximizing the likelihood [20]
@@
$=\sum_{i\leq j}\min(x_{ij},0)$ $-\ln(1+e^{-\left|x_{ij}\right|})-(1-a_{ij})x_{ij}$ (9)
subject to $\vec{b}_{ik}\geq 0,\vec{c}_{ik}\geq 0$.
@@ III.2 Multi-Scale Model
The natural way of modelling $\left\{\mathbf{A}^{(\ell)}:\ell\geq 0\right\}$ is deploying renormalization models that have to be self-consistent across scales and allowing for arbitrary partitions $\left\{\Omega_{\ell}\right\}_{\ell\geq 0}$. The multi-scale model [17] is the only one that takes into account all these properties (cfr. [3] for homogeneous partitions). In addition, by equipping the (global) multi-scale model with node embedding $\left\{\vec{x}_{i}\right\}_{i\in[1,N]}$, the connection probability for the new multi-scale model (MSM) reads
$p_{ij}:=p(\vec{x}_{i},\vec{x}_{j},w_{i})=\begin{cases}\makebox[51.96254pt][l]{$1-e^{-\langle\vec{x}_{i},\vec{x}_{j}\rangle}$}i\neq j\\ \makebox[51.96254pt][l]{$1-e^{-\frac{1}{2}\|\vec{x}_{i}\|^{2}-w_{i}}$}i=j\end{cases}$ (10)
subject to $\vec{x}_{ik}\geq 0,w_{i}\geq-\frac{1}{2}\|\vec{x}_{i}\|^{2}$ since the probability has to be bounded ($p_{ij}\in[0,1]$). In principle, one should impose $\langle\vec{x}_{i},\vec{x}_{j}\rangle\geq 0$, but this is equivalent to restricting every component to be non-negative, i.e. $\vec{x}_{ik}\geq 0$ (see Supp.Mat. II.6). The norm $\|\vec{x}_{i}\|:=\langle\vec{x}_{i},\vec{x}_{i}\rangle$ is the one induced by the scalar product. In general, $\vec{x}_{i}$ could be interpret as the propensity of $i$ to create a link with another node $j$. Theoretically, this role in encoded by its modulus (significance) and its relative angle (similarity) with the other embeddings. On the other hand, $w_{i}$ rules the self-loop of node $i$. A more detailed interpretantion of the node embeddings is lacking in the literature, but it is out of the scope for this work.
As for LPCA, $\left\{\vec{x}_{i}\right\}_{[0,N-1]}$ are obtained by maximizing the likelihood
$\mathcal{L}\left(\mathbf{X}|\mathbf{A}\right)$ $:=\sum_{i\leq j}a_{ij}\ln\left(p_{ij}\right)+\left(1-a_{ij}\right)\ln\left(1-p_{ij}\right)$ (11)
$=\sum_{i\leq j}a_{ij}\ln\left(1-e^{-\langle\vec{x}_{i},\vec{x}_{j}\rangle}\right)-\left(1-a_{ij}\right)\langle\vec{x}_{i},\vec{x}_{j}\rangle$ (12)
where $\textbf{X}:=\begin{bmatrix}-\vec{x}_{1}^{T}-\\ \vdots\\ -\vec{x}_{N}^{T}-\end{bmatrix}\in\mathbb{R}^{N\times D}$ is the vertical stacking of the embedding vectors, i.e. $\mathbf{X}_{i*}=\vec{x}_{i}$. We optimized only the vectors of the structural inequivalent nodes (see Supp.Mat. II.4). The loop parameters $\left\{w_{i}\right\}_{i\in[1,N]}$ are fixed to
$w_{i}=\begin{cases}\makebox[28.10083pt][l]{$-\frac{1}{2}\|\vec{x}_{i}\|^{2}$}\quad\textnormal{ if }p_{ii}=1\\ \makebox[28.10083pt][l]{$\to\infty$}\quad\textnormal{ if }p_{ii}=0\end{cases}$ (13)
after the training of the node embeddings. Indeed, for each node $i$, the likelihood of the self-loops $i$ displays only one term related to the presence ($a_{ii}=1$) or absence of the self-loop ($a_{ii}=0$). Therefore, by maximizing the likelihood the reached values for the loop parameters will be exactly the ones set above.
III.2.1 Self-Consistency and Renormalization
The MSM enforces the equality among the renormalized $\widetilde{\mathbf{P}}^{(\ell)}$ and the coarse-grained $\check{\mathbf{P}}^{(\ell)}$ probabilities [^3] The notation for the different probabilities uses the $\hat{}$ (hat) to refer to the P with fitted parameters. On the contrary, the $\;\check{}\;$ (check) reuses the 0-level parameters (the opposite of $\hat{\quad}$), and the $\widetilde{\quad}$ (tilde) resembles the “S” of “summed” for the renormalized probabilities. with the same functional form ([17], Supp.Mat. II), namely
$\begin{cases}\widetilde{\mathbf{P}}^{(\ell)}\stackrel{{\scriptstyle!}}{{=}}\check{\mathbf{P}}^{(\ell)}\\[4.30554pt] \mathbf{P}^{(\ell)}\stackrel{{\scriptstyle!}}{{=}}\mathbf{P}^{(m)}\qquad\forall\ell\geq 0,m\geq 0\end{cases}$ (14)
These two conditions are called self-consistency and scale-invariance. Here, we will use again the subscript $\ell=0$ to highlight that $\left\{\vec{x}_{i_{0}},w_{i_{0}}\right\}_{i\in[1,N_{0}]}$ are the fitted vectors at finest level 00, whereas $\left\{\vec{x}_{I},w_{I}\right\}_{I\in[1,N_{\ell}]}$ are the parameters at the level $\ell\geq 0$ where we set the block-nodes as $I:=\Omega_{0\to\ell}(i_{0})$. Therefore, by assuming that the block-node parameters are the sum of the lower level ones, i.e.
@@ III.2.1 Self-Consistency and Renormalization
Equation 14 are satisfied by means of Equation 10 (see section II). In this way, the MSM at level $\ell$ is recovered by inserting them in the Equation 10, namely
$p_{IJ}:=p(\vec{x}_{I},\vec{x}_{J},w_{I})=\begin{cases}\makebox[45.8509pt][l]{$1-e^{-\langle\vec{x}_{I},\vec{x}_{J}\rangle}$}I\neq J\\ \makebox[45.8509pt][l]{$1-e^{-\frac{1}{2}\|\vec{x}_{I}\|^{2}-w_{I}}$}I=J\end{cases}$ (16)
Thanks to the summation rule, the MSM has a lower computational complexity with respect to the single-scale models (see subsection II.7) as they need to be refitted at every scale.
In this essay, we took advantage of this property by fitting the embeddings vector from the observed network to recover all the coarser $\ell-parameters$ for every $\ell\geq 0$ by means of Equation 15. Finally, the probability $p_{IJ}$ among $\ell-communities$ was obtained by inserting the $\left\{\vec{x}_{I},w_{I}\right\}_{I\in[1,N_{\ell}]}$ in Equation 10.
IV Applications
@@ IV.1 ING Input-Output Network
ING Bank N.V. regularly reports the economic transactions of all ING clients for different years. We focused on the payments for the year 2022 between ING firms by removing the individual or non-Dutch clients, the flows of money sent/received by a non-ING account and the payments circulating inside a firm, i.e. self-payments. Since ING is the biggest bank in The Netherlands [21], this gave us the possibility of analyzing a major portion of the market.
More precisely, we chose the year 2022 both to ease the numerical calculations and to avoid skewed distributions by the aftermath of the COVID-19 pandemic. However, the procedure may be easily replicated for other time intervals, e.g. 3 years span, quarterly, $\cdots$
@@
At the firm-to-firm (f2f) resolution, the dataset is composed by $N_{f2f}\approx 3.4\cdot 10^{5}\text{ nodes},L_{f2f}\approx 4\cdot 10^{4}\text{links}$ which imply a density $\rho_{f2f}\approx 3.5\cdot 10^{-5}$. Therefore, the network is big and sparse. From this network, we aggregated the firms by NAICS (North American Industry Classification System) codes and set an edge among two sectors if there was at least one link among the firms of each community (see Equation 3). Then, we filtered out the “Public Administration” (92), “Finance and Insurance” (52), “Management of Companies and Enterprises”/“Holdings” (55) sectors to retain a production ION. Roughly, the reasons underneath the payments from/to the sectors 52-55-92 are not directly connected with a product/service. In particular, the “Public Administration” fluxes includes taxes and fees; the “Finance and Insurance”, money management, e.g. loans, that are not part of the production chain of any good; and the “Management of Companies and Enterprises”/“Holdings” are a collection of business entities, controlling stocks in other companies. Lastly, we mapped, for simplicity, the 6-digits NAICS codes to integers, i.e. $111110\to 0,111120\to 1,\cdots$.
@@
This is a first application of the MSM on the multiscale structure built from the ION. Therefore, we studied only the economic relationships among the sectors discarding the directionality and the amount of money of the link. In other words, each edge is reciprocated and binary. For example [^4] To be precise, the notation for the nodes $i,j$ refers to the observed microscopic nodes, namely $i:=i_{0},j:=j_{0}$, if $w_{ij}$ was the total amount of money sent from $i\textnormal{ to }j$ at level 00, we reciprocated the weight by setting $w_{ij}\to w^{{}^{\prime}}_{ij}:=\frac{w_{ij}+w_{ji}}{2}$ [5] and $a_{ij}=a_{ji}=1$ if $w^{{}^{\prime}}_{ij}>0$ and imposed $a_{ij}=a_{ji}=0\textnormal{ if }w^{{}^{\prime}}_{ij}=0$ to create a binary edge. In this way, a bidirectional link is created, i.e. $a_{ij}=1$ every time there was at least one directed flow among two sectors. The empirical ION is composed by $N_{0}=972$ sectors and $L_{0}=1.4\cdot 10^{5}$ links, which imply a density $\rho_{0}\approx 0.29\gg\rho_{f2f}$ by 4 order of magnitudes.
IV.1.1 Coarse-Graining The ION

### v2

@@ III Methodology
To capture the hierarchical structure, we used both the LogisticPCA [14] and the vector-based multi-scale model (MSM). We chose LPCA as the representative of the single-scale models, due to its straightforward yet effective derivation. In contrast, to the best of our knowledge, the MSM is the only model that explicitly allows for a renormalization rule applicable to arbitrary node partitions.
@@ III.1 Logistic PCA
LPCA [^2] The authors have historically called LPCA the directed logistic model, and no acronym has been assigned for its undirected counterpart [14]. In this work, we focus on the symmetric version, and refer to it simply as LPCA for easiness. [14] is a probabilistic model which aims to classify each edge as existing (class $1$) or non-existing (class 0). This is analogous to a standard “spam / non-spam” classification problem [13], but applied to links, which do not possess explicit features - hence, the need for node embeddings.
The probability of an existing edge between the nodes $i$ and $j$ is given by
@@
were introduced by the Chanpuriya et colleagues to capture the “homophily” and “heterophily” natures of each node, respectively (see Appendix A). By defining the non-negative matrices as $\textbf{B}:=\begin{bmatrix}-\vec{b}_{1}^{T}-\\ \vdots\\ -\vec{b}_{N}^{T}-\end{bmatrix}$ and similarly for $\mathbf{C}$, the scalar product among them becomes $\mathbf{X}:=\mathbf{B}\mathbf{B}^{T}-\mathbf{C}\mathbf{C}^{T}$, so that each element is defined by $x_{ij}:=\langle\vec{b}_{i},\vec{b}_{j}\rangle-\langle\vec{c}_{i},\vec{c}_{j}\rangle\in\mathbb{R}$. The node embeddings are obtained by maximizing the likelihood [23]
@@
$=\sum_{i<j}\min(x_{ij},0)$ $-\ln(1+e^{-\left|x_{ij}\right|})-(1-a_{ij})x_{ij}$ (9)
subject to non-negative entries of the vectors $\vec{b}_{ik}\geq 0,\vec{c}_{ik}\geq 0$. It is worth noting that the original likelihood [14] requires summing over all the entries of the adjacency matrix, which is equivalent of summing over the upper triangular elements included the diagonal since $\mathbf{A}$ is symmetric. However, since the network metrics considered in our study exclude the self-loops, we restricted the optimization to the off-diagonal elements of the adjacency matrix.
@@ III.2 Multi-Scale Model
A natural approach to modeling the multi-resolution adjacency matrices $\left\{\mathbf{A}^{(\ell)}:\ell\geq 0\right\}$ is by means of the multi-scale model (MSM). Specifically, this model is inherently scale-invariant and allows for arbitrary partitions $\left\{\Omega_{\ell}\right\}_{\ell\geq 0}$ of the microscopic nodes. In this work, we equip the (global) multi-scale model [17] with node embeddings $\left\{\vec{x}_{i}\right\}_{i\in[1,N]}$, yielding the following connection probability between nodes $i$ and $j$:
$p_{ij}:=\begin{cases}\makebox[67.67029pt][l]{$1-e^{-\langle\vec{x}_{i},\vec{x}_{j}\rangle}$}i\neq j\\ \makebox[67.67029pt][l]{$1-e^{-\frac{1}{2}\|\vec{x}_{i}\|^{2}-w_{i}}$}\end{cases}$ (10)
where $\|\vec{x}_{i}\|:=\langle\vec{x}_{i},\vec{x}_{i}\rangle$ is the Euclidean norm induced by the scalar product. Note that the probability is subject to $\vec{x}_{ik}\geq 0,w_{i}\geq-\frac{1}{2}\|\vec{x}_{i}\|^{2}$ to ensure it is non-negative. Although one might consider explicitly enforcing $\langle\vec{x}_{i},\vec{x}_{j}\rangle:=\left\|\vec{x}_{i}\right\|\left\|\vec{x}_{j}\right\|\theta_{ij}\geq 0$, this condition is naturally satisfied by requiring non-negative components for each vector, i.e. $\vec{x}_{ik}\geq 0$ (see subsection B.2). Roughly, $\vec{x}_{i}$ can be interpreted as the propensity of the node $i$ to establish a connection with another nodes. More specifically, the magnitude of the embedding can be seen as the node’s overall activity, whereas the angle $\theta_{ij}$ between embeddings indicates similarity between the nodes $i$ and $j$. The parameter $w_{i}$, instead, indicates the propensity of node $i$ of having a self-interaction. While a precise interpretation of these vectors remains an open question, such discussion falls outside the scope of this work.
The node embeddings are learned by maximizing the following likelihood
$\mathcal{L}\left(\mathbf{X}|\mathbf{A}\right)$ $:=\sum_{i<j}a_{ij}\ln\left(p_{ij}\right)+\left(1-a_{ij}\right)\ln\left(1-p_{ij}\right)$ (11)
$=\sum_{i<j}a_{ij}\ln\left(1-e^{-\langle\vec{x}_{i},\vec{x}_{j}\rangle}\right)-\left(1-a_{ij}\right)\langle\vec{x}_{i},\vec{x}_{j}\rangle$
where
$\textbf{X}:=\begin{bmatrix}-\vec{x}_{1}^{T}-\\ \vdots\\ -\vec{x}_{N_{irred}}^{T}-\end{bmatrix}\in\mathbb{R}^{N_{irred}\times D},$
and $N_{irred}$ is the number of irreducible structural inequivalent nodes (see subsection B.5). We restricted the full likelihood to the off-diagonal likelihood as done for the LPCA. If needed, the MSM allows to fix the self-loops values, after the likelihood maximization, according to the following rule
$w_{i}=\begin{cases}\makebox[31.6731pt][l]{$-\frac{1}{2}\|\vec{x}_{i}\|^{2}$}\quad\textnormal{ if }p_{ii}=1\\ \makebox[31.6731pt][l]{$\to\infty$}\quad\textnormal{ if }p_{ii}=0\end{cases}$ (12)
as explained in see subsection B.6.
III.3 Embedding Dimension
The selection of the optimal embedding dimension is still an active Research problem [24, 25, 26]. Here, we adopted the “Minimum Description Length” principle, approximated by the Bayesian Information Criteria (BIC) [25]. In the Supplemental Material [20], we report the BIC scores for the dimensions $D=1,2,8,16$ and, for completeness, also the AIC scores [24]. According to BIC, the optimal dimension depends on the resolution levels. For the Input-Output Network (ION), $D^{opt}_{\ell\geq 1}=1$ whereas $D^{opt}_{\ell=0}=2$. In contrast, for the WTW -a less complex network-, the minimum BIC is obtained for $D^{opt}_{\ell\geq 0}=1$ (see Supplemental Material). The fact that for lower levels $D^{opt}$ increases was somehow expected, since a larger number of nodes introduces more heterogeneity requiring a bigger embedding dimension to capture the observed network. However, since this is the first time the MSM is proposed, we report results for all embedding dimensions $D=1,2,8,16$ to provide a comprehensive evaluation.
III.4 Summed Models
In this section, we present the renormalization procedure used to obtain the scale-invariant versions of the MSM and the LPCA. As previously discussed, the MSM is inherently scale-invariant by design (see subsection III.2). On the other hand, LPCA requires additional modifications to enforce this property (cfr. subsection A.2).
III.4.1 Summed MSM
To start with, let us consider the MSM. Scale-invariance requires that the renormalized probability $\mathbf{P}_{sum}^{(\ell)}$ equals the coarse-grained probability $\mathbf{P}_{cog}^{(\ell)}$, and that the functional form of the model remains unchanged across resolution level ([17], Appendix B). Formally, the two conditions reads
$\begin{cases}\mathbf{P}_{sum}^{(\ell)}&\stackrel{{\scriptstyle!}}{{=}}\mathbf{P}_{cog}^{(\ell)}\\[4.30554pt] \mathbf{P}^{(\ell)}&\stackrel{{\scriptstyle!}}{{=}}\mathbf{P}^{(m)}\quad\forall\ell\geq 0,m\geq 0\end{cases}$ (13)
The fitted vectors at $\ell=0$ are denoted as $\left\{\vec{x}_{i_{0}},w_{i_{0}}\right\}_{i\in[1,N_{0}]}$, while $\left\{\vec{x}_{I},w_{I}\right\}_{I\in[1,N_{\ell}]}$ are the embeddings at the coarser level $\ell\geq 0$, where we defined the block-nodes as $I:=\Omega_{0\to\ell}(i_{0})$. Assuming that the block parameters are obtained by summing those at the micro level, namely
@@ III.4.1 Summed MSM
it holds that the scale-invariance conditions (Equation 13) are satisfied by the MSM probability function Equation 10 (see Appendix B). Thus, the MSM at level $\ell$ is recovered by substituting the summed parameters into Equation 10, yielding
$p_{IJ}:=p(\vec{x}_{I},\vec{x}_{J},w_{I})=\begin{cases}\makebox[62.14705pt][l]{$1-e^{-\langle\vec{x}_{I},\vec{x}_{J}\rangle}$}I\neq J\\ \makebox[62.14705pt][l]{$1-e^{-\frac{1}{2}\|\vec{x}_{I}\|^{2}-w_{I}}$}I=J.\end{cases}$ (15)
The summation rule also reduces the computational complexity compared to the models that require to be refitted at every scale, e.g. LPCA (see subsection B.8).
III.4.2 Summed LPCA
In contrast to the MSM, enforcing consistency across different scales in LPCA is less straightforward. The key issue is the absence of a renormalization rule that allows to solve the system Equation 13 using the logistic function (see subsection A.2). Additionally, if the vectors [^3] The fitted parameters, at a specific level $\ell$, are identified by the notation $\hat{x}_{i}$. Accordingly, $\hat{p}$ refers to the activation function p with fitted parameters $\hat{x}_{i}$. are re-fitted independently at each scale, one looses the dependence of higher level parameters with the lower ones. This effectively treats each level as if it was generated by a different underlying process. This contradicts the hierarchical structure of the network, as the coarse-grain (low-resolution) matrix is uniquely derived from the microscopic one. These results indicate that only the 0-vectors to construct a self-consistent, higher-scale LPCA. Moreover, that only the functional invariance (second equation in Equation 13) can be enforced in practice.
Concretely, we adopt the same scheme used in the MSM, as there is no specification of a renormalization rule to be exploited. Specifically, the coarse parameters are
where $I:=\mathbf{\Omega}_{0\to\ell}(i_{0})=(\mathbf{\Omega}_{\ell}\circ\,\cdots\,\circ\mathbf{\Omega}_{0})(i_{0})$ denotes the block-nodes at level $\ell$. To enforce functional invariance, these summed parameters are substituted in the LPCA activation function, yielding
This approach allows imposing functional invariance while preserving a coherent relationship between embeddings across different scales. Consequently, we can compare the expected values from the summed LPCA and MSM models against the aggregated empirical values.
IV Data
Level $\ell$ $N_{\ell}$ Reduced $N_{\ell}$ $L_{\ell}$ $\rho_{\ell}$
0 972 972 136887 0.29
1 647 647 93659 0.45
2 303 303 30861 0.67
3 87 67 3364 0.89
Table 1: Table for the ION datasets reporting, for each level $\ell$, the number of nodes ($N_{\ell}$), the reduced number of nodes, the total number of links and the density.
Level $\ell$ $N_{\ell}$ Reduced $N_{\ell}$ $L_{\ell}$ $\rho_{\ell}$
0 182 177 9993 0.61
1 152 147 7305 0.64
2 112 119 4822 0.67
3 92 88 2728 0.65
4 62 58 1332 0.70
5 32 24 400 0.80
Table 2: Table for the WTW datasets reporting, for each level $\ell$, the number of nodes ($N_{\ell}$), the reduced number of nodes, the total number of links and the density.
@@ IV.1 ING Input-Output Network
ING Bank N.V. regularly reports the economic transactions of all ING clients for different years. We focused on firm-to-firm payments in the year 2022 between ING bank accounts, filtering out the transactions involving individual or non-Dutch clients, flows to or from non-ING account and the internal transfers within the same firm, i.e. self-payments. Since ING is the largest bank in The Netherlands [27], the dataset covers a major portion of the national market. More precisely, we selected the year 2022 to simplify the numerical calculations and to avoid distortions in the data from the aftermath of the COVID-19 pandemic. However, the procedure may be easily extended for other time intervals, such as multi-year spans or quarterly snapshots.
@@
At the firm-to-firm (f2f) resolution, the dataset consists of $N_{f2f}\approx 3.4\cdot 10^{5}\text{ nodes},L_{f2f}\approx 4\cdot 10^{6}\text{ links}$, yielding a density $\rho_{f2f}\approx 3.5\cdot 10^{-5}$. This classifies the network as large and sparse. To analyze the data at a higher level, we aggregated the firms according to their NAICS (North American Industry Classification System) codes, and set an edge among two sectors if there was at least one payment among firms belonging to one of them (see Equation 3). Then, to focus on the production structure of this graph, we excluded sectors such as “Public Administration” (92), “Finance and Insurance” (52), “Management of Companies and Enterprises”/“Holdings” (55) as they are not directly related with a product/service. In particular, “Public Administration” fluxes include taxes and fees; “Finance and Insurance”, loans, that are not part of the supply chain of a product; and “Management of Companies and Enterprises”/“Holdings” include business entities that own shares in multiple firm. Lastly, we mapped, for simplicity, the 6-digits NAICS codes to integers, i.e. $111110\to 1,111120\to 2,\cdots$.
@@
Our work is the first application of the MSM to the multiscale structure derived from the ION. Therefore, we focused on the economic relationships among the sectors, discarding both the directionality and the volume of monetary flow. In this context, each edge is undirected (reciprocated) and binary. For example [^4] The $i,j$ refers to the observed 0-nodes, namely $i:=i_{0},j:=j_{0}$, if $w_{ij}$ refers to the total amount of money sent from $i\textnormal{ to }j$ at level 0, we reciprocated the transaction by setting $w_{ij}\to w^{{}^{\prime}}_{ij}:=\frac{w_{ij}+w_{ji}}{2}$ [5]. Furthermore, an edge was regarded as active ($a_{ij}=a_{ji}=1$) if $w^{{}^{\prime}}_{ij}>0$ or absent otherwise. This results in a symmetrical and binary edge, whenever there was at least one flow between two sectors. At this stage, the ION consists of $N_{0}=972$ sectors and $L_{0}=1.4\cdot 10^{5}$ links, yielding a network density of $\rho_{0}\approx 0.29$, roughly 4 order of magnitudes higher than $\rho_{f2f}$.
IV.2 Coarse-Graining The ION